Alright. So Math is definately not my thing. I always have trouble in math, no matter how hard I try I cannot do it. So lately I have been asking for help from my friends. When i get frustrated I say "Just gimme the answer!!" I think part of my problem is that I have become so mad that math is my only troublesome subject, that I have given up on it. We are now leaaring 'transforming formulas.' Ex.) l = a +(n - 1)d ; n ** You solve for n or something. I don't know. I hope someone can help.
Answer by stanbon(48569)

Yes, solve for "n", as follows:
1=a+(n-1)d
1=a+nd-d
1-a+d=nd
n=(1-a+d)/d
Cheers,
stan H.